# Semiparametric Skew-Elliptical Distributions For High-Dimensional Graphical Models

### v1

@@ Appendix A Proofs of elliptical skew-(S)KEPTIC estimators
Kendall’s tau is an unbiased estimator: $\mathbb{E}(\hat{\tau}_{ij})=\tau_{ij}$, where $\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$ is the population version.
$\mathbb{P}\left(\left|\hat{\tau}_{ij}-\tau_{ij}\right|>t\right)=$
$=\mathbb{P}\left(\left|2\sin\left(\frac{\pi}{2}A_{\hat{\tau}_{ij}}\right)-\sin\left(\frac{\pi}{2}\tau_{ij}\right)\right|>t\right)$
$\leq\mathbb{P}\left(\left|2A_{\hat{\tau}_{ij}}-\tau_{ij}\right|>\frac{2}{\pi}t\right)$
As in the Proof in Theorem 4.1 we have: $(I-\hat{D}^{2})<I$. This implies:
$\mathbb{P}\left(\sup_{ij}|\hat{S}^{\tau,\alpha}-\Sigma^{0}|>t\right)\leq\exp\left(-\frac{nt^{2}}{4\pi^{2}}\right)$ (29)
This result is obtained by exploiting the fact that we can write Kendall’s tau in terms of the U-statistic and using the Hoeffding inequality. ∎

### v2

@@ Appendix A Proofs of elliptical skew-(S)KEPTIC estimators
Similarly to the case of Spearman’s rho, we have that $\hat{S}^{\tau,\alpha}_{ij}=$ $2\sin\left(\frac{\pi}{2}A_{\hat{\tau}_{ij}}\right)$. We introduce $\Sigma^{\tau}$, the correlation matrix using Kendall’s tau statistic. We know that $A_{\tau}=L^{\top}ML$, where $M=I-D^{2}$. By Definition: $\Sigma^{\tau}=L^{\top}D_{\kappa}L$. This implies that:
$\left\lVert\widehat{D}_{\kappa}\right\rVert_{\max}\geq\left\lVert\widehat{M}\right\rVert_{\max},$
$\left\lVert L^{\top}\widehat{D}_{\kappa}L\right\rVert_{\max}\geq\left\lVert L^{\top}\widehat{M}L\right\rVert_{\max},$
$\left\lVert\widehat{\Sigma}^{\tau}\right\rVert_{\max}\geq\left\lVert A_{\widehat{\tau}}\right\rVert_{\max},$
$\left\lVert 2\sin\left(\frac{\pi}{2}\widehat{\Sigma}^{\tau}\right)\right\rVert_{\max}\geq\left\lVert 2\sin\left(\frac{\pi}{2}A_{\widehat{\tau}}\right)\right\rVert_{\max}.$
[…]
$E(\hat{\tau}_{ij})=\tau_{ij},$
where
$\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$
is the population version.
$\text{pr}\left(\left\lVert 2\sin\left(\frac{\pi}{2}\hat{\Sigma}^{\tau}\right)-\Sigma^{0}\right\rVert_{\max}>t\right)\quad$ $=\text{pr}\left(\max_{ij}\ \left|2\sin\left(\frac{\pi}{2}\hat{\tau}_{ij}\right)-\sin\left(\frac{\pi}{2}\tau_{ij}\right)\right|>t\right)$
$\leq\text{pr}\left(\max_{ij}\pi\left|\hat{\tau}_{ij}-\tau_{ij}\right|>t\right)\quad\text{(Lipshitz function)}$
$\leq\sum_{ij}\text{pr}\left(\left|\hat{\tau}_{ij}-\tau_{ij}\right|>\frac{t}{\pi}\right)\quad\text{(Boole's inequality)}$
$\leq p^{2}\exp\left(-\frac{nt^{2}}{2\pi^{2}}\right)\quad\text{(Hoeffding's inequality)}$
Setting
$\beta=p^{2}\exp\left(-\frac{nt^{2}}{2\pi^{2}}\right).$
And solving for $t$, we get:
$t=\sqrt{\frac{2\pi^{2}\log\left(\frac{2p^{2}}{\beta}\right)}{n}}.$
When $\beta=\frac{1}{p}$, we have $\sqrt{\frac{2\pi^{2}}{n}\log\left({p^{3}}\right)}.$
A.5 Proof of Theorem 5.3
We start by proving $\|\hat{L}-L\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$ and $\|\hat{M}-M\|_{\max}O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$. Kendall’s tau and Spearman’s rho statistics can be rewritten in terms of U-statistics, and their estimates of correlation matrix achieve $O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$ as rate of convergence. Under assumption (a), we know that $\Sigma^{0}$ has bounded eigenvalues, which implies:
$\|\hat{L}-L\|_{\max}\leq\tau_{2}\|\hat{\Sigma}-\Sigma^{0}\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right).$
Regarding the convergence of the estimated $\hat{M}$ matrix, we know that it has diagonal elements defined as: $\hat{m}_{kk}=1-\frac{2}{\pi}\frac{\hat{\alpha}_{kk}}{1+\hat{\alpha}_{kk}}$ where $\hat{\alpha}$ parameters are estimated via the EM algorithm introduced in Abe et al., (2021), for which we assume that the regulation conditions of Wu, (1983) hold. Then $\|\hat{M}-M\|_{\max}O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$. We find bounds for $\|\hat{\Omega}-\Omega\|_{\max}$.
$\|\hat{\Omega}-\Omega\|_{\max}$ $=\|2\sin\left(\frac{\pi}{2}(\hat{L}^{\top}\hat{M}\hat{L}^{\top})\right)^{-1}-2\sin\left(\frac{\pi}{2}({L}^{\top}{M}{L}^{\top})\right)^{-1}\|_{\max}$
$\leq\|(\hat{L}^{\top}\hat{M}\hat{L})^{-1}-({L}^{\top}{M}{L})^{-1}\|_{\max}$
$=\|(\hat{L}\hat{M}^{-1}\hat{L}^{\top})-({L}{M}^{-1}{L}^{\top})\|_{\max}$
$=\|(\hat{L}-L)\hat{M}^{-1}\hat{L}^{\top}+L(\hat{M}^{-1}-M^{-1})\hat{L}^{\top}+LM^{-1}(\hat{L}-L)\|_{\max}$
We define $A_{1}=(\hat{L}-L)\hat{M}^{-1}\hat{L}^{\top}$, $A_{2}=L(\hat{M}^{-1}-M^{-1})\hat{L}^{\top}$, and $A_{3}=LM^{-1}(\hat{L}^{\top}-L^{\top}_{0})$.
$\|A_{1}\|_{\max}=\|(\hat{L}-L)\hat{M}^{-1}\hat{L}^{\top}\|_{\max}\leq\|\hat{L}-L\|_{\max}\|\hat{M}^{-1}\|_{\max}\|\hat{L}^{\top}\|_{1}$
$\|\hat{L}-L\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right),\quad\|\hat{M}^{-1}\|_{\max}=O(1),\quad\|\hat{L}^{\top}\|_{1}=O(s).$
Where $s$ are the non-zero elements in the $L$ matrix.
$A_{2}=\|L(\hat{M}^{-1}-M^{-1})\hat{L}^{\top}\|_{\max}\leq\|L\|_{\infty}\|\hat{M}^{-1}-M^{-1}\|_{\max}\|\hat{L}^{\top}\|_{1}$
Where $\|\hat{M}^{-1}-M^{-1}\|_{\max}=\max_{j}\left|\frac{1}{\hat{M}_{jj}}-\frac{1}{M_{0,jj}}\right|=\left|\frac{\hat{M}_{jj}-M_{0,jj}}{\hat{M}_{jj}M_{0jj}}\right|$. This means that: $\|\hat{M}^{-1}-M^{-1}\||_{\max}=O_{P}(\|\hat{M}-M\|_{\max})=O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$.
$\|A_{3}\|_{\max}$ $=\|LM^{-1}(\hat{L}^{\top}-L^{\top}_{0})\|_{\max}$
$\leq\|L\|_{\infty}\|M^{-1}\|_{\infty}\|\hat{L}^{\top}-L^{\top}_{0}\|_{\max}=O_{P}\left(s\sqrt{\frac{\log p}{n}}\right)$
We conclude that:
$\|\hat{\Omega}-\Omega\|_{\max}=O_{P}\left(s\sqrt{\frac{\log p}{n}}\right).$
A.6 Proof of Theorem 5.4
As discussed in Lam and Fan, (2009), we only need to prove a bound in $\|\hat{\Omega}-\Omega\|_{\max}$ for the three shrinkage procedures proposed to recovery $\Omega$. For the glasso algorithm, the proof follows the proof of Theorem 2, in Lam and Fan, (2009). For the CLIME, the bounds are provided in Theorem 4 in Cai et al., (2011), under the Frobenius norm, which implies $\|\cdot\|_{\max}$. And for the graphical Dantzig selector, the bounds are provided in Section 3 of Yuan, (2010) work.